OpenStudy (anonymous):

solve the stem of equations 25x^2 + 9y^2 = 225 25x^2 - 9y^2 = 225

hartnn (hartnn):

subtract them , what u get ?

OpenStudy (anonymous):

18y^2 =450

OpenStudy (calculusfunctions):

450??? Really??

hartnn (hartnn):

actually, subtracting right side, u get 225-225 = ??

OpenStudy (anonymous):

ooo o ya lol soo 18y^2=0 y^2=18 y= 3radical 2

OpenStudy (calculusfunctions):

WHAT??

OpenStudy (calculusfunctions):

NO!

hartnn (hartnn):

18y^2 = 0 dividing 18 on both sides. whats 0/18 = ??

OpenStudy (anonymous):

oops ugh im sorry idk ive been studying too long.. its 0

hartnn (hartnn):

yes, so since y^2 = 0 y will also be = ... ?

OpenStudy (anonymous):

0

OpenStudy (calculusfunctions):

YEAH!!!

hartnn (hartnn):

now put y=0 in any equation and find x ...

OpenStudy (anonymous):

+or - 3

OpenStudy (calculusfunctions):

APPLAUSE!!!!

OpenStudy (anonymous):

lol

hartnn (hartnn):

good work :) you are \(\huge \color{red}{\checkmark}\)

OpenStudy (anonymous):

thank you

OpenStudy (calculusfunctions):

And a standing ovation for @hartnn

hartnn (hartnn):

LOL, THANK YOU SO MUCH SIR :)

OpenStudy (calculusfunctions):

LOL